OpenStudy (anonymous):

Amanda studied for 5 hours, which was one third as many hours as Raj. For how many hours did Raj study? What is the correct solution method for this?

OpenStudy (anonymous):

where are you options?

OpenStudy (anonymous):

A. t/3=5. Multiply both sides by 3. Raj studied for 15 hours. B. 3t=5. Divide both sides by 3. Raj studied for 3/5 hours. C. t+3=5. Subtract 3 by both sides. Raj studied for 2 hours. D. t-5=3. Add 5 to both sides. Raj studied for 8 hours.

OpenStudy (anonymous):

@julieanneheck

OpenStudy (anonymous):

u know this is basic 5th grader work

OpenStudy (anonymous):

right

OpenStudy (anonymous):

i remember learning it

OpenStudy (anonymous):

@rami2021 Listen pellethead, you forget a lot of pellet by the time you hit high school.

OpenStudy (anonymous):

um well did u try searching this online ts basic fractions

OpenStudy (anonymous):

I already got the answer

OpenStudy (anonymous):

ok